An electric toaster consumes 1875 W in operation. Assuming it is plugged into a 125 V source, what current flows in the circuit?
1
Expert's answer
2018-05-25T08:30:08-0400
P = 1875 W
U = 125 V
According to the electric power formula for the resistive load:
P = U*I → I=P/U
I= (1875 W)/(125 V)=15 A